### v1

@@ 4.4.1 Cost Analysis and Comparison
4.5 Ablation Study
4.5.1 Cost-Efficiency Analysis of Collage Sizes
@@
Table 7: Comparative Analysis of Random Initialization Approaches, Brute Force Solution, and our LCP algorithm for $2\times 2$ and $3\times 3$ Grid Collages.
@@
$K$ Method Steps Time Accuracy
$3\times 3$ Random Init. - - 28.1%
Brute Force 1,500 9.856 32.8%
LCP 500 4.479 33.7%
$2\times 2$ Random Init. - - 39.4%
Brute Force 24 0.164 45.1%
LCP 9 0.107 45.7%
Table 7: Comparative Analysis of Random Initialization Approaches, Brute Force Solution, and our LCP algorithm for $2\times 2$ and $3\times 3$ Grid Collages.
Figure 4: Average Prediction Accuracy by Position in 2x2 and 3x3 Grids.
4.5.2 Comparison of Optimization Methods
Figure 5: Case visualization.
4.6 Case Study
The position statistics depicted in Figure 4 reveal variations in prediction accuracy across different positions within $2\times 2$ and $3\times 3$ grids. Notably, images positioned in the top-left corner exhibit the highest probability of correct predictions, whereas those positioned in the bottom-left corner demonstrate the lowest accuracy. For example, a “room divider” is more accurately recognized when placed in the top-left, as opposed to being mislabeled as a “bed” in other positions. Similarly, “wombat” and “guacamole” samples, considered easier, maintain high accuracy regardless of placement, except when positioned in the least accurate bottom-left, which even led to a misprediction of the “guacamole.” This pattern is further exemplified by “letter opener” and “necklace” examples. These findings emphasize the importance of considering spatial arrangement when employing collage prompting in image recognition tasks, suggesting that optimal placement of images can substantially influence the GPT-4V model’s performance.
5 Conclusion
In conclusion, our study introduces Collage Prompting as a novel and cost-efficient method for leveraging GPT-4V in image recognition tasks, demonstrating that optimized image arrangements significantly enhance accuracy while reducing inference costs. Through extensive experimentation, we highlight the advantages of our approach over traditional methods, offering a practical solution to the high costs associated with large multi-modal models. This work not only advances the field of visual recognition but also sets the stage for future research into cost-effective inference strategies for advanced AI models.
Appendix 0.A Appendix
0.A.1 Experimental Details
0.A.1.1 Network Details.
The overall architecture of the collage predictor is depicted in Figure 6. It comprises multiple graph convolutional and pooling layers. The graph convolutional layers aggregate information from neighboring nodes, while the graph pooling layers retain the sub-graph information for each node. This structure preserves the basic graph structural information and facilitates message passing. By learning graph representations hierarchically and summarizing the node representations in each layer using readout functions, the graph representations are then input into a multi-layer perception (MLP) to perform graph regression prediction tasks, specifically predicting the overall accuracy of the collage graph.
Figure 6: Network architecture of collage predictor.
0.A.1.2 Details of Predictor Training.
For training the collage predictor, both evaluation datasets for $2\times 2$ and $3\times 3$ collage prompt are split into training and validation sets at a 9:1 ratio, where 90% of the data is allocated for training and the remaining 10% for validation. The collage predictor is trained with a batch size of 512 and a learning rate of 0.001 for 500 epochs. We utilize the Mean Squared Error (MSE) loss function during training. The node input feature dimension of the collage graph network is set to 512. The network architecture consists of three convolutional layers and pooling layers, with a pooling ratio of 0.5. We employ the Adam optimizer to optimize the model. The dimensions of the three MLP layers are set to [256, 128, 64]. During training, we utilize an early stopping strategy to prevent overfitting.
0.A.1.3 Details of LCP.
The pseudocode for our LCP algorithm is provided in Algorithm 1. When predicting arrangements using LCP, we employ uniform crossover without allowing duplicate genes and random mutation to introduce variation in the predicted arrangements for both $2\times 2$ and $3\times 3$ collage prompts. For the $3\times 3$ collage prompt, we initialize the population with 100 arrangements. In each generation, we select the top 20 arrangements with the highest accuracy to serve as parents for crossover and mutation. The evolution process continues for 10 generations, and we terminate it when the saturation threshold reaches 3. Similarly, for the $2\times 2$ collage prompt, we begin with an initial population of 5 arrangements. We retain the top 3 arrangements in each generation based on accuracy for further reproduction. The evolution process runs for 5 generations, and we stop it when the saturation threshold also reaches 3. Finally, we evaluate the predicted best arrangements by feeding them in batches of 4 to the GPT-4V API to obtain the actual prediction accuracy.
Figure 7: The process of Crossover and Mutation in the proposed LCP.
0.A.1.4 Details of Crossover and Mutation.
During the iterative process of optimizing arrangements using LCP, crossover and mutation of arrangements are involved. The specific processes of crossover and mutation are illustrated in Figure 7. At the initial stage of each iteration, our LCP algorithm predicts the accuracy of each initial arrangement using the collage predictor and retains the top-k collage arrangements. Then, any two arrangements from the top-k are randomly selected for node crossover to obtain n partial initial node arrangements for the collages. Finally, the remaining nodes are randomly allocated (mutated) to the blank positions in the collages.
0.A.2 Additional Visualized Results
@@
2 Initialise generation 0:;
3 $k:=0$;
4 $P_{k}$ := a set of $n$ randomly-generated arrangements;
5 Evaluate $P_{k}$:;
6 Compute $fitness(i)$ for each $i\in P_{k}$;
7 while not stop-criterion do

### v2

@@ 4.2 Baseline: Learning to Collage
By iteratively employing these steps, the initial arrangements are updated efficiently until the maximum iterations are achieved. After obtaining the arrangement with the best-expected accuracy, we can apply this arrangement configuration to enhance the performance of collage prompts. Algorithm 1 outlines the steps of the LCP algorithm, while Appendix C provides more details of the baseline method. To support reproducibility, we have released the dataset, baseline code, and model weights on our project page.
4.3 Metrics
To effectively assess the performance of our collage prompting approach, we utilize cost (i.e., $C_{n\times n}$) and accuracy (i.e., $A_{n\times n}$) as primary evaluation metrics, where $n\times n$ denotes the size of collage prompt. Collage prompt can significantly reduce costs but often at the expense of recognition performance loss. Besides evaluating cost and accuracy separately, we introduce two new metrics for a more comprehensive analysis as follow:
Cost-Effective Ratio (CER): CER applies logarithmic transformations on both primary metrics to improve the distinction and manageability of the values. It is formulated as,
$\mathrm{CER}=\frac{\left(\log(A_{n\times n}+1)\right)^{\gamma}}{\log(C_{n\times n}+e)},$ (2)
Precision-Cost Efficiency (PCE): PCE signifies the cost saved per accuracy loss, which is formulated as,
$\mathrm{PCE}=e^{\left(\frac{\lvert C_{n\times n}-C_{1\times 1}\rvert}{\lvert A_{n\times n}-A_{1\times 1}\rvert}\right)}.$ (3)
These two metrics balance the trade-offs between accuracy and cost, providing deeper insights into the efficiency of various collage configurations. We also use these metrics to evaluate the performance of our baseline algorithm.
5 Experiment
In this section, we benchmarked GPT-4V’s zero-shot collage prompt recognition performance on ImageNet-1K Russakovsky et al. (2015) and 10 other datasets (e.g., Caltech101 Fei-Fei et al. (2004), OxfordPets Parkhi et al. (2012), StanfordCars Krause et al. (2013), Flowers102 Nilsback and Zisserman (2008), Food101 Bossard et al. (2014), Aircraft Maji et al. (2013), SUN397 Xiao et al. (2010), DTD Cimpoi et al. (2014), EuroSAT Helber et al. (2019), UCF101 Soomro et al. (2012)). For the $1\times 1$ collage prompt, we referenced the zero-shot experiment results from GPT4Vis Wu et al. (2023c). Using the API service provided by OpenAI, we evaluated the recognition performance for $2\times 2$ and $3\times 3$ collage prompts. The specific model version used was “gpt-4-1106-vision-preview”. We used low-resolution to input images and set a random seed to ensure deterministic results.
5.1 Benchmark Results of Collage Prompting
Our analysis reveals that utilizing Collage Prompting with GPT-4V for image recognition significantly reduces inference costs without substantially compromising accuracy as shown in Table 1. While the ImageNet-1K dataset presents greater challenges due to long text labels, leading to a more drop in accuracy, accuracy on other medium-sized datasets remains substantial. By employing different configurations of collage prompting, including larger grid sizes ($2\times 2$ and $3\times 3$), we demonstrated a significant decrease in usage costs—approximately to 1/4 and 1/9 of the cost for single images, respectively. Despite a decrease in Top-1 accuracy as grid size increases, our baseline methods for optimizing these collage arrangements significantly reduce accuracy loss, achieving over 5% higher accuracy than random arrangements. The baseline approach highlights the balance between cost efficiency and performance preservation, making it a practical solution for leveraging large multi-modal models like GPT-4V in resource-constrained scenarios.
Our analysis demonstrates the effectiveness of collage prompting in enhancing cost-efficiency across various datasets, with $n\times n$ grid collages significantly outperforming single $1\times 1$ images in terms of the Cost-Effective Ratio (CER) as shown in Table 1. The $3\times 3$ grids, optimized through our collage graph optimization method, show the most notable improvements in cost efficiency, especially in datasets with simpler labels and less challenging images. Additionally, our Precision-Cost Efficiency (PCE) analysis underscores the trade-off between cost savings and accuracy loss, highlighting that our optimized $2\times 2$ and $3\times 3$ grid arrangements achieve substantial cost savings while minimizing accuracy loss as demonstrated in Table 1, thereby offering a balanced approach to cost-efficient image recognition with GPT-4V. Overall, these results underscore the practicality and efficiency of collage prompting in leveraging large multi-modal models for image recognition tasks under budget constraints.
5.2 Cost Analysis
We analyzed the costs associated with using AWS cloud servers for training and inference of traditional CNN and ViT models on the ImageNet-1k dataset, comparing the expenses with those of collage prompting using GPT-4V. Training ResNet-50 and ViT-B/16 from scratch incurred significant costs (\$2,091.63 and \$3,909.46, respectively). In contrast, leveraging GPT-4V for single $1\times 1$ image prediction reduced the cost to \$641.25. Additionally, collage prompting with $2\times 2$ or $3\times 3$ grid configurations further decreased costs to \$170.27 and \$87.75, respectively. The Cost-Effective Ratio (CER) highlights the stark contrast between collage prompting and traditional models. For example, the $3\times 3$ grid configuration has a CER 17 times higher than that of ViT-B/16, showcasing the cost efficiency of collage prompting.
Notably, while models like ViT and CLIP offer impressive capabilities, they entail certain barriers such as data, training, and computational resources. In contrast, GPT-4V enables zero-shot recognition with minimal setup, offering a significant advantage, especially when using collage prompting. This advantage extends to various visual recognition tasks, further emphasizing the practicality and efficiency of GPT-4V in real-world applications.
6 Limitation and Future work
Limitations. 1) Accuracy Drop: While collage prompting significantly reduces costs, it does trade off some recognition performance. Despite challenges with datasets like ImageNet-1K, accuracy on medium-sized datasets remains reasonable. Collage prompt is still applicable for tasks with lower accuracy requirements, such as image or video captioning. Our platform will help researchers enhance collage recognition performance, moving closer to the accuracy of standard prompting. 2) Collage Prompt in Other LLMs: We also tested collage prompting on other open-source and closed-source multimodal vision-language models (e.g., LLAVA-1.5, Gemini 1.5 Pro) and found that these models performed poorly in visual recognition tasks. These models generated non-existent or incorrect labels, produced repetitive outputs, and failed to recognize images within the collage prompt. Appendix F provides examples of these failures.
Future Work. In this paper, we propose a budget-friendly task of collage prompting for GPT-4V’s visual recognition and construct a benchmark for learning to optimize the collage prompt. Future work could explore text prompt optimization, visual prompting techniques to learn adversarial noise perturbations, LCP optimization for multiple arrangement candidates, and few-shot/many-shot methods to improve accuracy. Additionally, we will actively maintain and update CollagePrompt, expanding the baseline library, applying it to other multi-modal foundation models, and extending it to broader visual recognition tasks.
Acknowledgement
We acknowledge the National Computational Infrastructure of Australia for providing computational resources via the Sydney NCI Scheme. This work was supported in part by the Australian Research Council under Projects DP240101848 and FT230100549.
Appendix A Impact Statement
While collage prompting offers significant cost savings, especially for large-scale image recognition, it also introduces potential societal implications. The accuracy drop associated with collage prompting could have adverse effects in fields where precise recognition is paramount, such as medical imaging. These implications underscore the necessity for continued exploration into the reliability and safety of leveraging collage prompting for visual recognition tasks, particularly in fields where accuracy is critical.
A.1 Datasheet for CollagePrompt
Here, we provide a datasheet Gebru et al. (2021) for documenting and ensuring responsible usage of the CollagePrompt Benchmark.
1. Motivation
  • For what purpose was the dataset created? This dataset was created as a benchmark for studying the use of collage prompting to reduce the cost of GPT-4V while maintaining reasonable accuracy. It is intended for training and evaluating learning-based collage prompting optimization algorithms.
  • Who created the dataset (e.g., which team, research group) and on behalf of which entity (e.g., company, institution, organization)? The dataset was created by the authors of this paper.
  • Who funded the creation of the dataset? The creation of the dataset was funded by the Australian Research Council under Projects DP210101859 and FT230100549.
Composition
  • What do the instances that comprise the dataset represent (e.g., documents, photos, people, countries)? The dataset comprises the original information from CollagePrompt’s training and validation sets, as well as GPT-4V’s prediction results for the collage prompts.
  • How many instances are there in total (of each type, if appropriate)? This dataset includes over 110,000 2x2 and 100, 000 3x3 collage prompts with various arrangements of GPT-4V prediction results.
  • Does the dataset contain all possible instances or is it a sample (not necessarily random) of instances from a larger set? We randomly sampled 100,000 images from ImageNet to construct the 2x2 and 3x3 collage prompts.
  • What data does each instance consist of? Each instance consists of a collage and the corresponding GPT-4V prediction results.
  • Are relationships between individual instances made explicit? Different arrangements of collage prompts and their prediction results from the same set of images are grouped together.
  • Are there recommended data splits? Yes, the data splits for the training and validation sets are detailed in the JSON files.
  • Are there any errors, sources of noise, or redundancies in the dataset? The number of collage images may exceed the number of collage prompts prediction results. We have pre-cleaned the dataset to remove erroneous and unusable predictions of collage prompts.
  • Is the dataset self-contained, or does it link to or otherwise rely on external resources (e.g., websites, tweets, other datasets)? The dataset requires downloading the original ImageNet-1K dataset and the downstream evaluation datasets. Using the JSON files that record the original collage information, the code can construct the collage image dataset, which is then used along with our collage prompts prediction results.
  • Does the dataset contain data that might be considered confidential (e.g., data that is protected by legal privilege or by doctor-patient confidentiality, data that includes the content of individuals’ non-public communications)? No.
  • Does the dataset contain data that, if viewed directly, might be offensive, insulting, threatening, or might otherwise cause anxiety? No.
Collection Process
  • How was the data associated with each instance acquired? Each collage is constructed from the image datasets according to the order provided in the JSON file. The constructed collage, along with text prompts and dataset category labels, is then input into GPT-4V to obtain the prediction results for each sub-image in the collage.
  • What mechanisms or procedures were used to collect the data (e.g., hardware apparatuses or sensors, manual human curation, software programs, software APIs)? Each collage prompt is formatted and input into the GPT-4V API provided by OpenAI to obtain the corresponding prediction results. These results are then post-processed into a standardized, readable format.
  • Who was involved in the data collection process (e.g., students, crowd workers, contractors), and how were they compensated (e.g., how much were crowd workers paid)? The data collection process did not involve any manual labor; only API usage fees were incurred.
  • Over what timeframe was the data collected? The final version of the CollagePrompt dataset was collected in March 2024.
Uses
  • Has the dataset been used for any tasks already? Yes, we have used this dataset to train and evaluate our baseline algorithms for optimizing collage prompts.
  • Is there a repository that links to any or all papers or systems that use the dataset? Yes, https://collageprompting.github.io/.
Distribution
  • Will the dataset be distributed to third parties outside of the entity (e.g., company, institution, organization) on behalf of which the dataset was created? Yes, the dataset is publicly available online for anyone to access.
  • How will the dataset be distributed (e.g., tarball on website, API, GitHub)? The dataset can be downloaded on GitHub.
  • Will the dataset be distributed under a copyright or other intellectual property (IP) license, and/or under applicable terms of use (ToU)? Our dataset is distributed under CC BY 4.0. All codes on the GitHub repository are distributed under the MIT license.
  • Have any third parties imposed IP-based or other restrictions on the data associated with the instances? No.
  • Do any export controls or other regulatory restrictions apply to the dataset or to individual instances? No.
Maintenance
  • Who will be supporting/hosting/maintaining the dataset? The authors of this paper are supporting/maintaining the dataset.
  • Is there an erratum? No.
  • Will the dataset be updated (e.g., to correct labeling errors, add new instances, delete instances)? Please check the dataset web page or GitHub repository for any updates.
  • If others want to extend/augment/build on/contribute to the dataset, is there a mechanism for them to do so? Yes, they can use the code provided on our GitHub repository to generate data.
A.2 Data Hosting, Licensing, and Maintenance
The CollagePrompt Benchmark is licensed under the CC BY 4.4, and the data is hosted on Google Drive. All code used for data collection and developing baseline algorithms is distributed under the MIT license. The documentation and model checkpoints are also available on the GitHub repository. The Collage Prompt website (https://collageprompting.github.io/) is the central hub for all related information, including any future updates and maintenance.
Appendix B CollagePrompt Benchmark
B.1 Prompt Details
Our text prompts must include the collage filenames and category labels. After extensive experimentation, we have developed a stable version that ensures GPT-4V outputs the correct JSON format without any unrelated content. We also tested various prompting engineering techniques, but they did not significantly affect prediction accuracy. Our prompts can input single or multiple collages for prediction. When using the API, we set the batch size to 4, which does not significantly differ from predicting individual collages.
Text prompts used for $2\times 2$ collage
2x2_prompt = "I want you to act as an Image Classifier. I will provide you with few 2x2 grid collages and a list of optional categories. Your task is to choose the most relevant category for each of the nine images in the grid. Start with the top-left image of each grid and proceed left to right, then down each row. Assign a number index started with 0 for each image in the grid. Provide the prediction in a dict format for each grid collage, key is the number index, and value is the most relevant category for each image in the grid. The final output is also a dictionary. The key is image name of each grid collage, and the value is the prediction for each grid collage in a dict format. Do not provide explanations for your choices or any additional information just the dictionary of predictions in a JSON format. Only output the predictions in one JSON dictionary. Here is the image([]) and its optional categories([]). You have to choose strictly among the given categories and do not give any predictions that are not in the given category."
Text prompts used for $3\times 3$ collage
3x3_prompt = "I want you to act as an Image Classifier. I will provide you with few 3x3 grid collages and a list of optional categories. Your task is to choose the most relevant category for each of the nine images in the grid. Start with the top-left image of each grid and proceed left to right, then down each row. Assign a number index started with 0 for each image in the grid. Provide the prediction in a dict format for each grid collage, key is the number index, and value is the most relevant category for each image in the grid. The final output is also a dictionary. The key is image name of each grid collage, and the value is the prediction for each grid collage in a dict format. Do not provide explanations for your choices or any additional information just the dictionary of predictions in a JSON format. Only output the predictions in one JSON dictionary. Here is the image([]) and its optional categories([]). You have to choose strictly among the given categories and do not give any predictions that are not in the given category."
Datasets Classes Samples Label Tokens
ImageNet-1K 1,000 50,000 4,834
Caltech101 100 2,465 428
OxfordPets 37 3,669 203
StanfordCars 106 8,041 1,814
Flowers102 102 2,463 562
Food101 101 30,300 513
FGVCAircraft 100 3,333 565
SUN397 397 19,850 1,859
DTD 47 1,692 211
EuroSAT 10 8,100 49
UCF101 101 3,783 526
B.2 Evaluation Datasets
We evaluate the performance of collage prompt on ImageNet-1K and 10 other common downstream image recognition datasets. Table 4 presents statistics regarding the number of test samples and label tokens for each dataset. Label tokens represent the number of tokens encoded by the GPT-4 tokenizerhttps://platform.openai.com/tokenizer, providing a measure of the textual labels’ complexity for each dataset.
B.3 Dataset Format
Our dataset is structured as shown below. The training set begins with a random uniform sampling of images from the ImageNet-1k training dataset, followed by the construction of 2x2 and 3x3 collages according to the JSON files containing collage info in the metainfo directory. The evaluation datasets for all downstream datasets are formatted similarly to ImageNet-1k validation, constructing collages based on the standard data segmentation methods. All evaluation datasets consist of complete validation sets derived from ImageNet-1K.
[…]
Collage Information JSON
The format of all ‘collage_info.json’ files is consistent. Each file contains the original image name, category label, and position index within the collage for each sub-image. Using this JSON file and our provided code on GitHub, users can construct collage images and evaluate the prediction results of collage prompts from GPT-4V. The contents of a collage information JSON file are shown below:
a72bca2a3e.jpeg’: {
’0’: {’image’: ILSVRC2012_val_00024101.JPEG’,
synset_id’: 866,
label’: tractor’,
index’: 0},
’1’: {’image’: ILSVRC2012_val_00011876.JPEG’,
synset_id’: 480,
label’: automated teller machine’,
index’: 1},
’2’: {’image’: ILSVRC2012_val_00042300.JPEG’,
synset_id’: 842,
label’: swim trunks / shorts’,
index’: 2},
’3’: {’image’: ILSVRC2012_val_00034749.JPEG’,
synset_id’: 98,
label’: red-breasted merganser’,
index’: 3}},
d150b4fd58.jpeg’: {
’0’: {’image’: ILSVRC2012_val_00008159.JPEG’,
synset_id’: 776,
label’: saxophone’,
index’: 0},
’1’: {’image’: ILSVRC2012_val_00042315.JPEG’,
synset_id’: 123,
label’: spiny lobster’,
index’: 1},
’2’: {’image’: ILSVRC2012_val_00017726.JPEG’,
’3’: {’image’: ILSVRC2012_val_00016843.JPEG’,
synset_id’: 507,
label’: combination lock’,
index’: 3}},
Collage Prediction JSON. The prediction results of GPT-4V for collage prompts are preprocessed and stored in JSON format for ease of use, and then saved as Pickle files to conserve storage space. Files with the suffix ‘collage_pred_info.pkl’ contain the prediction results for each collage prompt, including the original image names and their positions within the collage. Below is an example of the contents of such a file:
{’0b73a3623d.jpeg’: {’ord’: [1, 2, 0, 3],
pred’: [0, 0, 1, 1],
ori’: [’n07697313_12937.JPEG’,
n03871628_38116.JPEG’,
n04493381_61391.JPEG’,
n02086910_6135.JPEG’]},
eb46c53c56.jpeg’: {’ord’: [3, 2, 1, 0],
pred’: [1, 1, 0, 1],
ori’: [’n07697313_12937.JPEG’,
n03871628_38116.JPEG’,
n04493381_61391.JPEG’,
n02086910_6135.JPEG’]},
’789e579a78.jpeg’: {’ord’: [3, 0, 2, 1],
pred’: [1, 1, 0, 0],
ori’: [’n07697313_12937.JPEG’,
n03871628_38116.JPEG’,
n04493381_61391.JPEG’,
n02086910_6135.JPEG’]},
f6810cfeb6.jpeg’: {’ord’: [0, 2, 3, 1],
pred’: [1, 0, 1, 0],
ori’: [’n07697313_12937.JPEG’,
n03871628_38116.JPEG’,
n04493381_61391.JPEG’,
n02086910_6135.JPEG’]},
’2ed874ee56.jpeg’: {’ord’: [3, 0, 1, 2],
ori’: [’n04335435_13775.JPEG’,
n04023962_6300.JPEG’,
n04507155_1825.JPEG’,
n04447861_2306.JPEG’]},
These files can be used to construct the collage image sets and to train and evaluate baseline methods for optimizing collage prompts. We also provide a complete download link for the collage image sets on GitHub, which can be used to reproduce the experimental results.
Figure 6: Network architecture of collage predictor.
Appendix C Experimental Details
Network Details. The overall architecture of the collage predictor is depicted in Figure 6. It comprises multiple graph convolutional and pooling layers. The graph convolutional layers aggregate information from neighboring nodes, while the graph pooling layers retain the sub-graph information for each node. This structure preserves the basic graph structural information and facilitates message passing. By learning graph representations hierarchically and summarizing the node representations in each layer using readout functions, the graph representations are then input into a multi-layer perception (MLP) to perform graph regression prediction tasks, specifically predicting the overall accuracy of the collage graph.
Details of Predictor Training. For training the collage predictor, both evaluation datasets for $2\times 2$ and $3\times 3$ collage prompt are split into training and validation sets at a 9:1 ratio, where 90% of the data is allocated for training and the remaining 10% for validation. The collage predictor is trained with a batch size of 512 and a learning rate of 0.001 for 500 epochs. We utilize the Mean Squared Error (MSE) loss function during training. The node input feature dimension of the collage graph network is set to 512. The network architecture consists of three convolutional layers and pooling layers, with a pooling ratio of 0.5. We employ the Adam optimizer to optimize the model. The dimensions of the three MLP layers are set to [256, 128, 64]. During training, we utilize an early stopping strategy to prevent overfitting. We trained the model for approximately 8 hours using an Nvidia GPU RTX 2080TI.
Details of LCP. The pseudocode for our LCP algorithm is provided in Algorithm 1. When predicting arrangements using LCP, we employ uniform crossover without allowing duplicate genes and random mutation to introduce variation in the predicted arrangements for both $2\times 2$ and $3\times 3$ collage prompts. For the $3\times 3$ collage prompt, we initialize the population with 100 arrangements. In each generation, we select the top 20 arrangements with the highest accuracy to serve as parents for crossover and mutation. The evolution process continues for 10 generations, and we terminate it when the saturation threshold reaches 3. Similarly, for the $2\times 2$ collage prompt, we begin with an initial population of 5 arrangements. We retain the top 3 arrangements in each generation based on accuracy for further reproduction. The evolution process runs for 5 generations, and we stop it when the saturation threshold also reaches 3. Finally, we evaluate the predicted best arrangements by feeding them in batches of 4 to the GPT-4V API to obtain the actual prediction accuracy.
Figure 7: The process of Crossover and Mutation in the proposed LCP.
@@ Appendix C Experimental Details
2
3Initialise generation 0:;
4
[…]
6
[…]
8
9Evaluate $P_{k}$:;
10
11Compute $fitness(i)$ for each $i\in P_{k}$;
12
13while not stop-criterion do
14